A 51.9-kg person, running horizontally with a velocity of +3.43 m/s, jumps onto a 14.2-kg sled that is initially at rest. (a) Ignoring the effects of friction during the collision, find the velocity of the sled and person as they move away. (b) The sled and person coast 30.0 m on level snow before coming to rest. What is the coefficient of kinetic friction between the sled and the snow?
Answer #1

a) Accoridng to the conservation of momentum

          m1 u 1 + m2 u 2 = ( m1 + m2 ) V

initially sled is at rest. so u2 = 0

          V = m1 u1 / ( m1 + m2)

               = (51.9 kg) (3.43 m/s) / ( 51.9 kg + 14.2 kg )

               = 2.69 m/s

b)

Work done is equal to change in kinetic energy

    ( 1/2) m v2 = ( mu) mg d

coefficient of kinetic friction

           mu = v2 / 2 g d

                 = ( 2.69 m/s) 2 / 2 ( 9.8 m/s2) (30.0 m)

                 = 0.0123